OpenStudy (anonymous):

Medal What is 66 2/3% of 327?

OpenStudy (acxbox22):

use a calculator or notice that 66 2/3% =2/3 so 2/3 x 327 = ??

OpenStudy (anonymous):

what do you mean

OpenStudy (acxbox22):